The coefficient of x50 in (1 + x)41 (1 - x + x2)40 is
 
  1
 
  2
 
  3
 
 

Grade:12th Pass

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear Vishwesh

(1 + x)41 (1 - x + x2)40

=(1+x)(1 + x)40 (1 - x + x2)40

=(1+x)(1+x3)40

=(1+x3)40     +x(1+x3)40


both the term does not contain x50 term because 50 is not multiple of 3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin